OpenStudy (anonymous):

15 runners are competing in a race. if the top 3 runners are going to the olympics, how many different groups of 3 could be chosen for the olympics?

OpenStudy (anonymous):

you are being asked to compute the number of ways to choose 3 out of 15, written as \[\dbinom{15}{3}\] read "fifteen choose 3" and computed via \[\dbinom{15}{3}=\frac{15\times 14\times 13}{3\times 2}\] divide first, multiply last

OpenStudy (anonymous):

is that a combination? :)

jimthompson5910 (jim_thompson5910):

If order matters (ie if there's a first, second and third place), then you use a permutation If order doesn't matter (as long as you place in the top 3, who cares), then use a combination

jimthompson5910 (jim_thompson5910):

My guess is that there is no first, second or third distinctions for qualifying for the olympics, so order doesn't matter

OpenStudy (anonymous):

thank you both ! :D

jimthompson5910 (jim_thompson5910):

sure thing